# Learning Enhanced Ensemble Filters

### v1

@@ 7 Acknowledgements
Appendix A Supplementary Information
This appendix provides supplementary information supporting our Measure Neural Mapping enhanced Ensemble Filter (MNMEF) architecture presented in Section 4.
A.1 Pseudo Codes for Our Architecture
@@
Recall that in both the initial training and the fine-tuning process, we always work under Data Assumption 1: we fix dynamic operator $\Psi$ (time step $\Delta t$ implicitly embedded), observation operator $h$, and noise covariance matrices $\Sigma$ and $\Gamma$ and generate multiple trajectories of the state-observation system through independent driving noise realization sequences.
The core of our framework (Algorithm 1) is to update the ensemble from $\{\widehat{v}^{(n)}_{j}\}_{n=1}^{N}$ to $\{\widehat{v}^{(n)}_{j+1}\}_{n=1}^{N}$. This is a basic component for both the training and inference. Based on this, we can calculate the loss and do the initial training with the ensemble size $N$ according to Algorithm 2. When we need to proceed with a different ensemble size $N^{\prime}\neq N$, we follow Algorithm 3 to efficiently fine-tune part of the parameters.
@@
We first load the pre-trained parameter $\theta_{\mathrm{ST}}^{(N)}$ that remain unchanged regardless of ensemble size, along with the fine-tuned parameters $\theta_{\mathrm{gain}}^{(N^{\prime})}$, $\theta_{\mathrm{infl}}^{(N^{\prime})}$, and $\theta_{\mathrm{loc}}^{(N^{\prime})}$ that were specifically optimized for ensemble size $N^{\prime}$. We then initialize the ensemble $\{v_{0}^{(n)}\}_{n=1}^{N^{\prime}}$ according to equation (4.38). For each time step $j$, we evolve the ensemble from $\{v_{j}^{(n)}\}_{n=1}^{N^{\prime}}$ to $\{v_{j+1}^{(n)}\}_{n=1}^{N^{\prime}}$ from using Algorithm 1. This process can be extended to observation trajectories of arbitrary length. At each time step, we can approximate the true state $v_{j}^{\dagger}$ using the ensemble mean $\bar{v}_{j}=\frac{1}{N^{\prime}}\sum_{n=1}^{N^{\prime}}v_{j}^{(n)}$.
@@
1:  Get predicted states from $\{v_{j}^{(n)}\}_{n=1}^{N}$ to $\{\widehat{v}^{(n)}_{j+1}\}_{n=1}^{N}$ (4.39a).
2:  Get predicted observations from $\{\widehat{v}^{(n)}_{j+1}\}_{n=1}^{N}$ to $\{\widehat{y}^{(n)}_{j+1}\}_{n=1}^{N}$(4.39b).
@@
6:  Process with neural networks for inflation and localization:
[…]
@@
$v^{(n)}_{j+1}=(\widehat{v}^{(n)}_{j+1}+\widehat{u}^{(n)}_{\theta})+(K_{\theta})_{j+1}\left(y^{\dagger}_{j+1}-\widehat{y}^{(n)}_{j+1}\right)$
10:  return  Ensemble $\{v^{(n)}_{j+1}\}_{n=1}^{N}$.
@@
0:  Dataset hyperparameters $J,M$; ensemble size $N$.
@@
A.2 Multihead Attention
@@
$\mathsf{A}(u,w)(j)=\frac{\sum_{j=1}^{M}\exp\bigl{(}\langle Qu(j),Kw(k)\rangle\bigr{)}Vw(k)}{\sum_{\ell=1}^{M}\exp\bigl{(}\langle Qu(j),Kw(\ell)\rangle\bigr{)}},\quad\forall j\in[N].$ (A.1)
@@
For each head $r$, we compute the attention output $\mathsf{A}_{r}(u,w)$ according to (A.1). The multihead attention mechanism $\mathsf{A}^{R}(u,w)$ combines the outputs from $R$ heads through concatenation followed by linear projection. For $j\in[N]$,
$\mathsf{A}^{R}(u,w)(j)=W^{O}\bigl{[}\mathsf{A}_{1}(u,w)(j);\mathsf{A}_{2}(u,w)(j);\ldots;\mathsf{A}_{R}(u,w)(j)\bigr{]},$ (A.2)
@@
$\theta(\mathsf{A}^{R})=\{Q_{r},K_{r},V_{r}\}_{r=1}^{R}\cup\{W^{O}\}.$ (A.3)
In Section 3, we view attention as an operator on probability measures. This formulation naturally extends to multihead attention. The multihead attention operator acts on probability measures similarly to standard attention, but utilizes $R$ parallel attention heads. In addition, we introduce two major properties of the set transformer architecture in Subsection 4.1. These two properties are preserved for multihead attention since multihead attention with $R$ heads effectively performs $R$ standard attention operations in parallel and concatenates their results.
Appendix B Implementation Details
@@
B.1 An Example of Learning the Localization Weight Matrix
$d_{\mathrm{peri}}(k,\ell)=\min\bigl{(}\lvert k-\ell\rvert,40-\lvert k-\ell\rvert\bigr{)}.$ (B.1)
Then we can consider all the unique distance values given by:
$\{D_{k\ell}=d_{\mathrm{peri}}(k,\ell)|k,\ell\in[40]\}=\{0,1,\ldots,20\}.$ (B.2)
$K=\widehat{C}^{vh}\circ L^{vh}\left(\widehat{C}^{hh}\circ L^{hh}+\Gamma\right)^{-1},$ (B.3)
we have
$\begin{split}(L^{vh})_{k,\ell}&=g(d_{\mathrm{peri}}(k,4\ell)),\quad k=1,2,\ldots,40,\quad\ell=1,2,\ldots,10,\\ (L^{hh})_{k,\ell}&=g(d_{\mathrm{peri}}(4k,4\ell)),\quad k=1,2,\ldots,10,\quad\ell=1,2,\ldots,10.\end{split}$ (B.4)
In our learning framework introduced in Subsection 4.2.3, we learn a vector $\widehat{g}\in\mathbb{R}^{21}$ according to (4.37) for the function values $g_{\theta}(D)$ for $D=0,1,2,\ldots,21$. Then in our proposed scheme (4.35),
$K_{\theta}=K_{\theta}^{(1)}\circ L_{\theta}^{(1)}\left(K_{\theta}^{(2)}\circ L_{\theta}^{(2)}+\Gamma\right)^{-1},$ (B.5)
we can calculate the parameterized localization weight matrices by
$\begin{split}(L_{\theta}^{(1)})_{k,\ell}&=g_{\theta}(d_{\mathrm{peri}}(k,4\ell)),\quad k=1,2,\ldots,40,\quad\ell=1,2,\ldots,10,\\ (L_{\theta}^{(2)})_{k,\ell}&=g_{\theta}(d_{\mathrm{peri}}(4k,4\ell)),\quad k=1,2,\ldots,10,\quad\ell=1,2,\ldots,10.\end{split}$ (B.6)
B.2 Feature Dimensions
@@
$\mathsf{F}^{\textrm{ST}}(u)=\mathsf{F}^{\textrm{NN}}_{2}(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})\circ\mathsf{F}^{\textrm{Cat}}(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})\circ\mathsf{S}^{\textrm{ST}}_{4}(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})\circ\mathsf{S}^{\textrm{ST}}_{3}(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})\circ\mathsf{C}^{\textrm{ST}}(s,\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})\circ\mathsf{S}^{\textrm{ST}}_{2}(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})\circ\mathsf{S}^{\textrm{ST}}_{1}(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})\circ\mathsf{F}^{\textrm{NN}}_{1}(u),$ (B.7)
where $\mathsf{F}^{\textrm{NN}}_{1}$ and $\mathsf{F}^{\textrm{NN}}_{2}$ are multilayer perceptrons consisting of multiple feedforward layers and activation layers (3.8); $\mathsf{S}^{\textrm{ST}}_{1,2}$ and $\mathsf{S}^{\textrm{ST}}_{3,4}$ are SABs serving as the encoder and decoder respectively; $\mathsf{F}^{\textrm{Cat}}$ is a layer that concatenate all elements in a sequence into a long feature vector. Note that layers $\mathsf{S}^{\textrm{ST}}_{1}$, $\mathsf{S}^{\textrm{ST}}_{2}$, $\mathsf{C}^{\textrm{ST}}$, $\mathsf{S}^{\textrm{ST}}_{3}$ and $\mathsf{S}^{\textrm{ST}}_{4}$ contain attention mechanisms, which are implemented as multihead attention with 8 heads (A.2).
@@
B.3 Hyperparameter Optimization via Grid Search
For each of the five benchmark methods mentioned above, we perform grid search on the inflation parameter $\alpha>1$. After the analysis step, the state $v^{(n)}$ is updated according to:
$v^{(n)}\leftarrow v^{(n)}+(\alpha-1)\left(v^{(n)}-\frac{1}{N}\sum_{\ell=1}^{N}v^{(\ell)}\right),\quad n=1,2,\ldots,N.$ (B.8)
In addition to inflation, for LETKF specifically, we consider the localization radius parameter $r$, which is used in the Gaspari–Cohn (GC) function [45] $\mathrm{GC}(d;\sqrt{10/3}r)$ to map distances to localization weights.

### v2

@@ 7 Acknowledgements
Appendix A Overview of the Architecture and our Complete Learning Framework
@@
Recall that in both the initial training and the fine-tuning process, we always work under Data Assumption 1: we fix dynamic operator $\Psi$ (time step $\Delta t$ implicitly embedded), observation operator $h$, and noise covariance matrices $\Sigma$ and $\Gamma$. We generate multiple trajectories of the state-observation system through independent realizations of the driving noise sequences.
The core of our framework (Algorithm 1) is to update the ensemble from $\{v^{(n)}_{j}\}_{n=1}^{N}$ to $\{v^{(n)}_{j+1}\}_{n=1}^{N}$. This is a basic component for both the training and inference. Based on this, we can calculate the loss and do the initial training with the ensemble size $N$ according to Algorithm 2. When we need to proceed with a different ensemble size $N^{\prime}\neq N$, we follow Algorithm 3 to efficiently fine-tune part of the parameters.
@@
We first load the pre-trained parameter $\theta_{\mathrm{ST}}^{(N)}$ that remain unchanged regardless of ensemble size, along with the fine-tuned parameters $\theta_{\mathrm{gain}}^{(N^{\prime})}$, $\theta_{\mathrm{infl}}^{(N^{\prime})}$, and $\theta_{\mathrm{loc}}^{(N^{\prime})}$ that were specifically optimized for ensemble size $N^{\prime}$. We then initialize the ensemble $\{v_{0}^{(n)}\}_{n=1}^{N^{\prime}}$ according to equation (4.37). For each time step $j$, we evolve the ensemble from $\{v_{j}^{(n)}\}_{n=1}^{N^{\prime}}$ to $\{v_{j+1}^{(n)}\}_{n=1}^{N^{\prime}}$ using Algorithm 1. This process can be extended to observation trajectories of arbitrary length. At each time step, we can approximate the true state $v_{j}^{\dagger}$ using the ensemble mean $\overline{v}_{j}=\frac{1}{N^{\prime}}\sum_{n=1}^{N^{\prime}}v_{j}^{(n)}$.
@@
1:  Get predicted states $\{\widehat{v}^{(n)}_{j+1}\}_{n=1}^{N}$ from $\{v_{j}^{(n)}\}_{n=1}^{N}$ (4.38a).
2:  Get predicted observations $\{\widehat{y}^{(n)}_{j+1}\}_{n=1}^{N}$ from $\{\widehat{v}^{(n)}_{j+1}\}_{n=1}^{N}$ (4.38b).
@@
6:  Process with the neural network for localization:
@@
$v^{(n)}_{j+1}=\widehat{v}^{(n)}_{j+1}+(K_{\theta})_{j+1}\left(y^{\dagger}_{j+1}-\widehat{y}^{(n)}_{j+1}\right)$
10:  Process with the neural network for inflation:
[…]
11:  Update the estimation with the learned inflation term $\widehat{u}^{(n)}_{\theta}$ by
$v^{(n)}_{j+1}\leftarrow v^{(n)}_{j+1}+\widehat{u}^{(n)}_{\theta}.$
12:  return  Ensemble $\{v^{(n)}_{j+1}\}_{n=1}^{N}$.
@@
0:  $J,M\in\mathbb{N}$; ensemble size $N$.
@@
Appendix B Description of Multihead Attention
@@
$\mathsf{A}(u,w)(j)=\frac{\sum_{j=1}^{M}\exp\bigl{(}\langle Qu(j),Kw(k)\rangle\bigr{)}Vw(k)}{\sum_{\ell=1}^{M}\exp\bigl{(}\langle Qu(j),Kw(\ell)\rangle\bigr{)}},\quad\forall j\in[N].$ (B.1)
@@
For each head $r$, we compute the attention output $\mathsf{A}_{r}(u,w)$ according to (B.1). The multihead attention mechanism $\mathsf{A}^{R}(u,w)$ combines the outputs from $R$ heads through concatenation followed by linear projection. For $j\in[N]$, we compute the output as
$\mathsf{A}^{R}(u,w)(j)=W^{O}\bigl{[}\mathsf{A}_{1}(u,w)(j);\mathsf{A}_{2}(u,w)(j);\ldots;\mathsf{A}_{R}(u,w)(j)\bigr{]},$ (B.2)
@@
$\theta(\mathsf{A}^{R})=\{Q_{r},K_{r},V_{r}\}_{r=1}^{R}\cup\{W^{O}\}.$ (B.3)
In Section 3, we view attention as an operator on probability measures. This formulation naturally extends to multihead attention. The multihead attention operator acts on probability measures similarly to standard attention, but utilizes $R$ parallel attention heads. In addition, we introduce two major properties of the set transformer architecture in Subsection 4.1, i.e. the permutation invariance and the adaptation to variable input lengths. These two properties are preserved for the version using multihead attention, since multihead attention with $R$ heads effectively performs $R$ standard attention operations in parallel and concatenates their results.
Appendix C Implementation Details
@@
C.1 Feature Dimensions
@@
$\mathsf{F}^{\textrm{ST}}(u)=\mathsf{F}^{\textrm{NN}}_{2}(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})\circ\mathsf{F}^{\textrm{Cat}}(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})\circ\mathsf{S}^{\textrm{ST}}_{4}(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})\circ\mathsf{S}^{\textrm{ST}}_{3}(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})\circ\mathsf{C}^{\textrm{ST}}(s,\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})\circ\mathsf{S}^{\textrm{ST}}_{2}(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})\circ\mathsf{S}^{\textrm{ST}}_{1}(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})\circ\mathsf{F}^{\textrm{NN}}_{1}(u),$ (C.1)
where $\mathsf{F}^{\textrm{NN}}_{1}$ and $\mathsf{F}^{\textrm{NN}}_{2}$ are multilayer perceptrons consisting of multiple feedforward layers and activation layers (3.8); $\mathsf{S}^{\textrm{ST}}_{1,2}$ and $\mathsf{S}^{\textrm{ST}}_{3,4}$ are SABs serving as the encoder and decoder respectively; $\mathsf{F}^{\textrm{Cat}}$ is a layer that concatenate all elements in a sequence into a long feature vector. Note that layers $\mathsf{S}^{\textrm{ST}}_{1}$, $\mathsf{S}^{\textrm{ST}}_{2}$, $\mathsf{C}^{\textrm{ST}}$, $\mathsf{S}^{\textrm{ST}}_{3}$ and $\mathsf{S}^{\textrm{ST}}_{4}$ contain attention mechanisms, which are implemented as multihead attention with 8 heads (B).
@@
C.2 An Example of Learning the Localization Weight Matrix
$d_{\mathrm{peri}}(k,\ell)=\min\bigl{(}\lvert k-\ell\rvert,40-\lvert k-\ell\rvert\bigr{)}.$ (C.2)
Then we can consider all the unique distance values given by:
$\{D_{k\ell}=d_{\mathrm{peri}}(k,\ell)|k,\ell\in[40]\}=\{0,1,\ldots,20\}.$ (C.3)
$K=\widehat{C}^{vh}\circ L^{vh}\left(\widehat{C}^{hh}\circ L^{hh}+\Gamma\right)^{-1},$ (C.4)
we have
$\begin{split}(L^{vh})_{k,\ell}&=g(d_{\mathrm{peri}}(k,4\ell)),\quad k=1,2,\ldots,40,\quad\ell=1,2,\ldots,10,\\ (L^{hh})_{k,\ell}&=g(d_{\mathrm{peri}}(4k,4\ell)),\quad k=1,2,\ldots,10,\quad\ell=1,2,\ldots,10.\end{split}$ (C.5)
In our learning framework introduced in Subsection 4.2.3, we learn a vector $\widehat{g}\in\mathbb{R}^{21}$ according to (4.36) for the function values $g_{\theta}(D)$ for $D=0,1,2,\ldots,20$. Then in our proposed scheme (4.34),
$K_{\theta}=K_{\theta}^{(1)}\circ L_{\theta}^{(1)}\left(K_{\theta}^{(2)}\circ L_{\theta}^{(2)}+\Gamma\right)^{-1},$ (C.6)
we can calculate the parameterized localization weight matrices by
$\begin{split}(L_{\theta}^{(1)})_{k,\ell}&=g_{\theta}(d_{\mathrm{peri}}(k,4\ell)),\quad k=1,2,\ldots,40,\quad\ell=1,2,\ldots,10,\\ (L_{\theta}^{(2)})_{k,\ell}&=g_{\theta}(d_{\mathrm{peri}}(4k,4\ell)),\quad k=1,2,\ldots,10,\quad\ell=1,2,\ldots,10.\end{split}$ (C.7)
C.3 Initial Conditions
Here we provide details of the initial conditions for the experiments in Section 5. To ensure our initial states lie on the attractor, we first select a base value and then evolve the system’s dynamics forward for a substantial number of time steps. This serves as a “burn-in” period. The initial conditions detailed below for each system are subsequently used to generate all trajectories for pretraining, fine-tuning, and testing.
  1. 1.
    Lorenz ’96 (Subsection 5.2): Sample $x_{0}\sim\mathsf{N}(5,I_{40})$ and run the forward dynamic (5.5) for random $10^{3}$ to $5\times 10^{5}$ time steps with step size $\Delta t=0.15$ to get $v^{\dagger}_{0}$. Within each $\Delta t$ time interval, there are $5$ RK4 numerical integration steps with the step size $\Delta t/5=0.03$. The initial distribution is set to be $\mathsf{N}(v^{\dagger}_{0},I_{40})$.
  2. 2.
    Kuramoto–-Sivashinsky (Subsection 5.3): Set the interval $[0,L]$ as $L=32\pi$. For $u_{0}(x)=\cos(2x/L)(1+\sin(2x/L))$, run the forward dynamic (5.6) for random $10^{3}$ to $5\times 10^{5}$ time steps with step size $\Delta t=1.00$ to get $\widehat{u}$. Within each $\Delta t$ time interval, there are $4$ RK4 numerical integration steps with the step size $\Delta t/4=0.25$. The initial state $v^{\dagger}_{0}=(\widehat{u}(x_{0}),\ldots,\widehat{u}(x_{127}))$ with $x_{\ell}=\ell L/128$, $\ell=0,1,\ldots,127$. The initial distribution is set to be $\mathsf{N}(v^{\dagger}_{0},I_{128})$.
  3. 3.
    Lorenz ’63 (Subsection 5.4): Sample $x_{0}\sim\mathsf{N}(0,I_{3})$ and run the forward dynamic (5.7) for random $10^{3}$ to $5\times 10^{5}$ time steps with step size $\Delta t=0.15$ to get $v^{\dagger}_{0}$. Within each $\Delta t$ time interval, there are $5$ RK4 numerical integration steps with the step size $\Delta t/5=0.03$. The initial distribution is set to be $\mathsf{N}(v^{\dagger}_{0},I_{3})$.
C.4 Hyperparameter Setting for Our Training
We provide details of the training hyperparameters in experiments. The parameters listed in Table 8 are utilized for pretraining with an ensemble size of $N=10$ in the experiments detailed in Section 5. When fine-tuning with a different ensemble size $N^{\prime}\neq N$ (Subsection 4.4), the number of training trajectories is halved compared to the quantity used for pretraining. Furthermore, the learning rate for fine-tuning is set to $1/10$ of the learning rate used during pretraining. We use the AdamW optimizer [54] for both the pretraining and fine-tuning.
During the training process, we implement a clamping mechanism to ensure numerical stability and prevent issues such as exploding estimated trajectories. If the absolute value of any dimension of an estimated particle exceeds the specified clamp value (provided in the “Clamp” column of Table 8), its sign is preserved, but its absolute value is replaced by the clamp threshold.
Table 8: Training hyperparameters for pretraining with an ensemble size N=10. For fine-tuning with N’ $\neq$ N, the number of training trajectories is halved, and the learning rate is reduced to $1/10$ of the values listed. Obs $\Delta t$ refers to the observation time step.
Dataset Train Traj Test Traj Obs Learning Batch Clamp
Num Length Num Length $\Delta t$ Rate Size
Lorenz ’96 8192 60 64 1500 0.15 1e-3 512 20
KS 8192 60 64 2000 1.00 5e-4 256 10
Lorenz ’63 8192 60 64 1500 0.15 1e-3 1024 60
C.5 Hyperparameter Optimization via Grid Search